Question 1
You want to prove that a population mean is not 7. Data is ratio and a large sample size will be used (n = 225). When the sample is done the sample mean is computed to be 7.2 and the sample standard deviation is 1.5. Test the hypothesis using an alpha of 5%. Use formal hypothesis testing and compute the p-value.
Question 2
An associate asks for your help in deciphering the following Anova table, indicating that a test of means was performed on the average age of people at three different locations. Unfortunately the person leaves before you can get any further explanation or detail, but you know the results must be presented before senior staff within twenty minutes!
Use the formal 5-step hypothesis testing model to describe the test, the results and interpret the results in conversational English.

ANOVA
Source of Variation SS df MS F P-value F crit
Between Groups 700.0 2 350.0 4.00 0.0288 3.32
Within Groups 2625.0 30 87.5

Total 3325.0 32

Answer by stanbon(29608)

You want to prove that a population mean is not 7. Data is ratio and a large sample size will be used (n = 225). When the sample is done the sample mean is computed to be 7.2 and the sample standard deviation is 1.5. Test the hypothesis using an alpha of 5%. Use formal hypothesis testing and compute the p-value.
Ho: u = 7
Ha: u is not equal to 7 (Claim)
------------------
test statistic: t(7.2) = (7.2-7)/[1.5/sqrt(225)] = 2
p-value: 2*P(t>2 with df= 224) = 0.0467
-----------
Since the p-value is less than 5%, reject Ho.
=====================================================

Question 2
An associate asks for your help in deciphering the following Anova table, indicating that a test of means was performed on the average age of people at three different locations. Unfortunately the person leaves before you can get any further explanation or detail, but you know the results must be presented before senior staff within twenty minutes!
Use the formal 5-step hypothesis testing model to describe the test, the results and interpret the results in conversational English.
---
Ho: The means are all equal
Ha: At least one of the means is not equal to the others.
--------------
Critical value for F = 3.32
Test statistic is F = 4
p-value = 0.0288
---------------
Conclusion: Since the p-value is less than 5%, reject Ho
=================================================================
Cheers,
Stan H.
ANOVA
Source of Variation SS.. df MS.... F. P-value F crit
Between Groups.... 700.0 2 350.0 4.00 0.0288.. 3.32
Within Groups.... 2625.0 30 87.5
Total............ 3325.0 32